# Automating Data Annotation under Strategic Human Agents: Risks and Potential Solutions

### v1

@@ 1 Introduction
Compared to prior studies on strategic classification under sequential settings [5, 6, 7] that mainly focused on developing a classifier robust to strategic agents, we study the long-term impacts of model retraining with agent strategic feedback. Instead of assuming actual labels are available while retraining, we consider more practical scenarios with model-annotated samples. Although the risk of using model-annotated samples to retrain models has been highlighted in existing works [3], ours is the first to incorporate strategic feedback from human agents. In App. C, we discuss more related works in detail. Our contributions are summarized as follows:
@@ 3.2 Social welfare: dynamics of qualification rate
Let $F_{X}(x)$ be the cumulative density function corresponding to $P_{X}$. Denote $\mathcal{J}=\{x|f_{0}(x)=0\}$ as the half-space in $\mathbb{R}^{d}$ determined by the classifier $f_{0}$ trained with $\mathcal{S}_{o,0}$. Under the retraining process, if $F_{X}$ and $P_{Y|X}(1|x)$ are convex on $\mathcal{J}$, then $q_{t+1}<q_{t}$, $\forall t\geq 1$.
Note that $q_{t+1}<q_{t}$ in Thm. 3.5 holds only for $t\geq 1$. Because agent behavior can only improve their labels, prior-best-response $q_{0}$ always serves as the lower bound of $q_{t}$. The half-space $\mathcal{J}$ in Thm. 3.5 specifies the region in feature space where agents have incentives to change their features. The convexity of $F_{X}$ and $P_{Y|X}(1|x)$ ensure that as $f_{t}$ evolves from $t=1$: (i) fewer agents choose to improve their features, and (ii) agents expect less improvement from feature changes. Thus, $q_{t}$ decreases over time. Conditions in Thm. 3.5 can be satisfied by common distributions $P_{X}$ (e.g., Uniform, $\text{Beta}(\alpha,\beta)$ with $\alpha>\beta$) and labeling functions $P_{Y|X}(1|x)$ (e.g., linear function, quadratic functions with degree greater than 1). The proof and a more general analysis are shown in App. G.5. We also show that Thm. 3.5 is valid under diverse experimental settings (Sec. 5, App. E, App. F).
@@ 3.4 Intervention to stabilize the dynamics
Specifically, at each time $t$, instead of adding $\mathcal{S}_{m-1,o}=\{x_{t-1}^{i},f_{t-1}(x_{t-1}^{i})\}_{i=1}^{N}$ (samples annotated by the model $f_{t-1}$) to training data $\mathcal{S}_{t}$ ((2)), we use the probabilistic model $h_{t-1}(x)$ to annotate each sample according to the following: For each sample $x$, we label it as $1$ with probability $h_{t-1}(x)$, and as 0 otherwise. Here $h_{t-1}(x)\approx D^{t-1}_{Y|X}(1|x)$ is the estimated posterior probability learned from $\mathcal{S}_{t-1}$ (e.g., using logistic model). We call the procedure refined retraining process if model-annotated samples are generated in this way based on a probabilistic sampler.
@@ 4.2 Impact of short-term fairness intervention
(i) Without fairness intervention, $\forall\sigma_{t}$, always there exists feature distributions and cost matrices under which group $j$ switches to be advantaged at $t+1$
@@ 5 Experiments
We conduct experiments on two synthetic (Uniform, Gaussian), one semi-synthetic (German Credit [18]), and one real dataset (Credit Approval [19]) to validate the dynamics of $a_{t},q_{t},\Delta_{t}$ and the unfairness. Note that only the Uniform dataset satisfies all assumptions and the conditions in our theoretical analysis, while the Gaussian and German Credit datasets violate the conditions in Thm. 3.5. The Credit Approval dataset violates all assumptions and conditions of the main paper. The decision-maker trains logistic regression models for all experiments using stochastic gradient descent (SGD) over $T$ steps. We present the experimental results of the Gaussian and German Credit datasets to illustrate the dynamics of $a_{t},q_{t},\Delta_{t}$ in this section, while the results for Uniform and Credit Approval data are similar and shown in App. E.
@@
German Credit dataset [18]. This dataset includes features for predicting individuals’ credit risks. It has 1000 samples and 19 numeric features, which are used to construct a larger-scaled dataset. Specifically, we fit a kernel density estimator for all 19 features to generate 19-dimensional features, the corresponding labels are sampled from the distribution $P_{Y|X}$ which is estimated from data by fitting a logistic classifier with 19 features. Given this dataset, the first 10 features are used to train the classifiers. The attribute "sex" is regarded as the sensitive attribute. The systematic bias is created by increasing/decreasing $P_{Y|X}$ by $0.05$. Other parameters $n,r,T,q_{0}$ are the same as Table 1. Since $P_{Y|X}$ is a logistic function, Assumption 3.2 can be satisfied easily as illustrated in App. F.1.
@@
Additional experiments. We provide more results in App. F to: (i) verify Thm. 3.3 and Thm. 3.6; (ii) Visualize the influence of each factor including training rounds, cost matrices, the ratio between human-annotated and model-annotated samples, and whether the agents are strategic; (iii) visualize the evolution of unfairness when different retraining strategies are applied to different groups.
@@ C.1 Strategic Classification
Strategic classification with label changes. Several other lines of literature enable strategic behaviors to cause label changes. The first line of literature mainly focuses on incentivizing improvement actions where agents have budgets to invest in different actions and only some of them cause the label change (improvement) [10, 32, 33, 34, 35, 36, 37, 38, 12]. The other line of literature focuses on causal strategic learning [39, 40, 31, 41, 42]. These works argue that every strategic learning problem has a non-trivial causal structure which can be explained by a structural causal model, where intervening on causal nodes causes improvement and intervening on non-causal nodes means manipulation.
@@ C.2 Bias amplification during retraining
There has been an extensive line of study on the computer vision field about how machine learning models amplify the dataset bias, while most works only focus on the one-shot setting where the machine learning model itself amplifies the bias between different groups in one training/testing round [46, 47]; Another work theoretically studies dynamic benchmarking [48] to propose a retraining scheme to increase model accuracy with adversarial human-annotation. In recent years, another line of research focuses on the amplification of dataset bias under model-annotated data where ML models label new samples on their own and add them back to retrain themselves[49, 50, 51, 52, 4, 53, 54, 55, 56]. These works study the bias amplification in different practical fields including resource allocation [54], computer vision [51], natural language processing [52] and clinical trials [56]. The most related work is [3] which studied the influence of retraining in the non-strategic setting. Also, there is a work [4] touching on the data feedback loop under performative setting, but it focused on empirical experiments under medical settings where the feature distribution shifts are mainly caused by treatment and the true labels in historical data are highly accessible. Besides, the data feedback loop is also related to recommendation systems. where extensive works have studied how the system can shape users’ preferences and disengage the minority population [57, 58, 59, 55]. However, previous literature did not touch on the retraining process.
@@ F.1 Additional Experimental Setups
Dynamics under different $r$. Although experiments in Sec. 5 and App. E already demonstrate the validity of Thm. 3.3, Prop. 3.4 and Thm. 3.6, the ratio $r$ of human-annotated examples is subject to change in reality. Therefore, we first provide results for $r\in\{0,0.05,0.1,0.3\}$ in all 3 datasets. $r$ only has small values since human-annotated examples are likely to be expensive to acquire. Fig. 15 shows all results under different $r$ values. Specifically, Fig. 15(a) and 15(b) show results for synthetic dataset 1, Fig. 15(c) and 15(d) show results for synthetic dataset 2, while Fig. 15(e) and 15(f) show results for German Credit data. On every row, $r=0.3,0.1,0.05,0$ from the left to the right. All figures demonstrate the robustness of the theoretical results, where $a_{t}$ always increases and $q_{t}$ decreases starting from $t=1$. $\Delta_{t}$ also has different dynamics as specified in Thm. 3.6.
@@ F.4 Additional Results on Fairness
In this section, we provide additional results on the dynamics of unfairness and classifier bias under different $r$ (the same settings in App. F.2). We aim to illustrate situation (iii) in Thm. 4.1 where the refined retraining process is applied on the advantaged group $i$ while the original process is applied on the disadvantaged group $j$. From Fig. 17(a), 17(c) and 17(e), we can see unfairness reaches a minimum in the middle of the retraining process, suggesting the earlier stopping of retraining brings benefits. From Fig. 17(b), 17(d) and 17(f), we can see $\Delta_{t}$ for the disadvantaged group $j$ reaches a minimum in the middle of the retraining process, but generally not at the same time when unfairness reaches a minimum.
@@ G.2 Derivation of factors influencing $a_{t},q_{t}$
According to the proof of Theorem 3.3, all agents with feature vector $x$ who are admitted by $f_{0}$ will be admitted by $f_{t}$ ($t>0$), making all agents who possibly improve must belong to $\mathcal{J}$. Since $F_{X}$ and $P_{Y|X}(1|x)$ are both convex and non-decreasing in $\mathcal{J}$, $P_{X}(x),U(x)$ are non-decreasing in $\mathcal{J}$. Then note that $f_{t}$ always lies below $f_{t-1}$, therefore, for each agent who improves at $t$ having feature vector $x_{i,t}$, we can find an agent at $t-1$ with corresponding $x_{i,t-1}$ such that both $P_{X}(x_{i,t-1})>P_{X}(x_{i,t})$ and $U(x_{i,t-1})>U(x_{i,t})$. This will ensure $I(t)<I(t-1)$. Thus, $q_{t}$ decreases starting from 1. ∎
Generally, (7) reveals that if $P_{X}(x)\cdot U(x)$ is convex in $\mathcal{J}$,$q_{t}$ decreases. This illustrates Prop. 3.5 gives a sufficient condition which is much easier to verify.
@@
At round $t$, we know $A_{\theta_{t}^{i}}^{it}>A_{\theta_{t}^{j}}^{jt}$. To reach demographic parity, the acceptance rates need to be the same, so at least one of the following situations must happen: (i) $\widetilde{\theta}_{t}^{i}>\theta_{t}^{i}$ and $\widetilde{\theta}_{t}^{j}>\theta_{t}^{j}$; (ii) $\widetilde{\theta}_{t}^{i}<\theta_{t}^{i}$ and $\widetilde{\theta}_{t}^{j}<\theta_{t}^{j}$; (iii) $\widetilde{\theta}_{t}^{i}<\theta_{t}^{i}$ and $\widetilde{\theta}_{t}^{j}<\theta_{t}^{j}$. Next we prove that (i) and (ii) cannot be true by contradiction.
@@
Since we want to sacrifice the least accuracy to reach demographic parity. Since the agent best response will never make them worse off, we will know $a_{t}^{i}$ corresponding to the acceptance rate of group $i$ under the optimal threshold $\theta_{t}^{i}$ is always larger than or equal to $a_{0}^{i}$. Thus, it suffices to show that under the condition of Thm. 4.2, $a_{t}^{j}<a_{0}^{i}$.
@@ Appendix H All plots with error bars
Although all our theoretical results are expressed in terms of expectation, we provide error bars for all plots in the main paper, App. E and F if randomness is applicable. All following figures demonstrate expectations as well as error bars ($\pm$ 1 standard deviation). Overall, the experiments have reasonable standard errors. However, experiments in the Credit Approval dataset [19] (Fig. 33) incur larger standard errors, which is not surprising because the dataset violates several assumptions. Finally, note that we conduct 50-100 randomized trials for every experiment and we should expect much lower standard errors if the numbers of trials become large.

### v2

@@ 1 Introduction
Compared to prior studies on strategic classification under sequential settings [5, 6, 7] that mainly focused on developing a classifier robust to strategic agents, we study the long-term impacts of model retraining with agent strategic feedback. Instead of assuming actual labels are available while retraining, we consider more practical scenarios with model-annotated samples. Although the risks on accuracy [3] and fairness [8] of using model-annotated samples to retrain models have been highlighted, ours is the first to incorporate strategic feedback from human agents. In App. C, we discuss more related works in detail. Our contributions are summarized as follows:
@@ 3.2 Social welfare: dynamics of qualification rate
Consider the setting where the $d$-dimensional feature space $X\in\mathbb{R}^{d}$ where $F_{X}(x),P_{X}(x),P_{Y|X}(1|x)$ are the cumulative distribution function, probability density function and the labeling when $Y=1$. Denote $\mathcal{J}=\{x|f_{0}(x)=0\}$ as the half-space in $\mathbb{R}^{d}$ determined by the classifier $f_{0}$. Under the retraining process, $\forall t\geq 1,~{}q_{t+1}\leq q_{t}$ if either of the following conditions holds: (i) $F_{X}$ and $P_{Y|X}(1|x)$ are convex on $\mathcal{J}$; (ii) for each dimension $x[i],i\in[d]$, $F_{X}(x)$ and $P_{Y|X}(1|x)$ are convex with respect to $x[i]$.
Note that $q_{t+1}\leq q_{t}$ in Thm. 3.5 holds only for $t\geq 1$. Because agent behavior can only improve their labels, prior-best-response $q_{0}$ always serves as the lower bound of $q_{t}$. The half-space $\mathcal{J}$ in Thm. 3.5 specifies the region in feature space where agents have incentives to change their features. The convexity of $F_{X}$ and $P_{Y|X}(1|x)$ ensure that as $f_{t}$ evolves from $t=1$: (i) fewer agents choose to improve their features, and (ii) agents expect less improvement from feature changes. Thus, $q_{t}$ decreases over time. Conditions in Thm. 3.5 can be satisfied by common distributions $P_{X}$ (e.g., Uniform, $\text{Beta}(\alpha,1)$ with $\alpha>1$) and labeling functions $P_{Y|X}(1|x)$ (e.g., linear function, quadratic functions with degree greater than 1). The proof and a more general analysis are shown in App. G.5. We also show that Thm. 3.5 is valid under diverse experimental settings (Sec. 5, App. E, App. F).
@@ 3.4 Intervention to stabilize the dynamics
Specifically, at each time $t$, instead of adding $\mathcal{S}_{m-1,o}=\{x_{t-1}^{i},f_{t-1}(x_{t-1}^{i})\}_{i=1}^{N}$ (samples annotated by the model $f_{t-1}$) to training data $\mathcal{S}_{t}$ (2), we use the probabilistic model $h_{t-1}(x)$ to annotate each sample according to the following: For each sample $x$, we label it as $1$ with probability $h_{t-1}(x)$, and as 0 otherwise. Here $h_{t-1}(x)\approx D^{t-1}_{Y|X}(1|x)$ is the estimated posterior probability learned from $\mathcal{S}_{t-1}$ (e.g., using logistic model). We call the procedure refined retraining process if model-annotated samples are generated in this way based on a probabilistic sampler.
@@ 4.2 Impact of short-term fairness intervention
(i) Without fairness intervention, $\forall\sigma_{t}$, there always exists feature distributions and cost matrices under which group $j$ switches to be advantaged at $t+1$;
@@ 5 Experiments
We conduct experiments on two synthetic (Uniform, Gaussian), one semi-synthetic (German Credit [19]), and one real dataset (Credit Approval [20]) to validate the dynamics of $a_{t},q_{t},\Delta_{t}$ and the unfairness  [^2] https://github.com/osu-srml/Automating-Data-Annotation-under-Strategic-Human-Agents. Note that only the Uniform dataset satisfies all assumptions and the conditions in our theoretical analysis, while the Gaussian and German Credit datasets violate the conditions in Thm. 3.5. The Credit Approval dataset violates all assumptions and conditions of the main paper. The decision-maker trains logistic regression models for all experiments using stochastic gradient descent (SGD) over $T$ steps. We present the experimental results of the Gaussian and German Credit datasets to illustrate the dynamics of $a_{t},q_{t},\Delta_{t}$ in this section, while the results for Uniform and Credit Approval data are similar and shown in App. E.
@@
German Credit dataset [19]. This dataset includes features for predicting individuals’ credit risks. It has 1000 samples and 19 numeric features, which are used to construct a larger-scaled dataset. Specifically, we fit a kernel density estimator for all 19 features to generate 19-dimensional features, the corresponding labels are sampled from the distribution $P_{Y|X}$ which is estimated from data by fitting a logistic classifier with 19 features. Given this dataset, the first 10 features are used to train the classifiers. The attribute "sex" is regarded as the sensitive attribute. The systematic bias is created by increasing/decreasing $P_{Y|X}$ by $0.06$. Other parameters $n,r,T,q_{0}$ are the same as Table 1. Since $P_{Y|X}$ is a logistic function, Assumption 3.2 can be satisfied easily as illustrated in App. F.1.
@@
Additional experiments. We provide more results in App. F to: (i) verify Thm. 3.3 and Thm. 3.6; (ii) Visualize the influence of each factor including training rounds, cost matrices, the ratio between human-annotated and model-annotated samples, whether the agents are strategic, and whether certain assumptions are violated; (iii) visualize the evolution of unfairness when different retraining strategies are applied to different groups.
@@ 6 Conclusion & Limitations
Acknowledgement
This material is based upon work supported by the U.S. National Science Foundation under award IIS-2202699 and IIS-2416895, by OSU President’s Research Excellence Accelerator Grant, and grants from the Ohio State University’s Translational Data Analytics Institute and College of Engineering Strategic Research Initiative.
@@ C.1 Strategic Classification
Strategic classification with label changes. Several other lines of literature enable strategic behaviors to cause label changes. The first line of literature mainly focuses on incentivizing improvement actions where agents have budgets to invest in different actions and only some of them cause the label change (improvement) [11, 33, 34, 35, 36, 37, 38, 39, 13, 40, 41]. The other line of literature focuses on causal strategic learning [42, 43, 32, 44, 45]. These works argue that every strategic learning problem has a non-trivial causal structure which can be explained by a structural causal model, where intervening on causal nodes causes improvement and intervening on non-causal nodes means manipulation.
@@ C.2 Bias amplification during retraining
There has been an extensive line of study on the computer vision field about how machine learning models amplify the dataset bias, while most works only focus on the one-shot setting where the machine learning model itself amplifies the bias between different groups in one training/testing round [49, 50]; Another work theoretically studies dynamic benchmarking [51] to propose a retraining scheme to increase model accuracy with adversarial human-annotation. In recent years, another line of research focuses on the amplification of dataset bias under model-annotated data where ML models label new samples on their own and add them back to retrain themselves[52, 53, 54, 55, 4, 56, 57, 58, 59]. These works study the bias amplification in different practical fields including resource allocation [57], computer vision [54], natural language processing [55], generative models [60, 61] and clinical trials [59]. The most related work is [3] which studied the influence of retraining in the non-strategic setting. Also, there is a work [4] touching on the data feedback loop under performative setting, but it focused on empirical experiments under medical settings where the feature distribution shifts are mainly caused by treatment and the true labels in historical data are highly accessible. Besides, the data feedback loop is also related to recommendation systems. where extensive works have studied how the system can shape users’ preferences and disengage the minority population [62, 63, 64, 58]. However, previous literature did not touch on the retraining process.
@@ D.4 Additional Discussions on refined retraining process
D.5 Additional Discussions on the linearity of the model
Our model is built upon previous works on strategic classification (e.g., [32, 12, 2]) where the decision policy was assumed to be transparent and interpretable since human agents expect to face a linear classifier which they can understand, especially under high-stake situations such as job hiring, college admission and loan application. Meanwhile, as pointed out by Raab and Liu [13], Zhang et al. [12], Levanon and Rosenfeld [67], in practical circumstances, the decision-maker can first fit a non-linear model to learn the embeddings of agents’ preliminary features and then produce the embedded features. The decision-maker then uses a linear classifier on the new set of features and agents also best respond with respect to new features. Under this setting, nonlinearity is involved but our results still hold. Practical examples include (i) FICO credit score [12]: FICO credit score is based on a complex set of features, and the decision-maker (e.g., a bank) simply uses a threshold of the score to assign decisions; (ii) Spam classification [67]: The decision-maker classifies spam using a linear classifier, but the features are produced as an embedding by a neural network based on number of words in the post, number of phone numbers in the post and number of followers of the user.
@@ F.1 Additional Experimental Setups
Dynamics under different $r$. Although experiments in Sec. 5 and App. E already demonstrate the validity of Thm. 3.3, Prop. 3.4 and Thm. 3.6, the ratio $r$ of human-annotated examples is subject to change in reality. Therefore, we first provide results for $r\in\{0,0.05,0.1,0.3\}$ in all 3 datasets. $r$ only has small values since human-annotated examples are likely to be expensive to acquire. Fig. 17 shows all results under different $r$ values. Specifically, Fig. 17(a) and 17(b) show results for synthetic dataset 1, Fig. 17(c) and 17(d) show results for synthetic dataset 2, while Fig. 17(e) and 17(f) show results for German Credit data. On every row, $r=0.3,0.1,0.05,0$ from the left to the right. All figures demonstrate the robustness of the theoretical results, where $a_{t}$ always increases and $q_{t}$ decreases starting from $t=1$. $\Delta_{t}$ also has different dynamics as specified in Thm. 3.6.
@@ F.4 Additional Results on Fairness
In this section, we provide additional results on the dynamics of unfairness and classifier bias under different $r$ (the same settings in App. F.2). We aim to illustrate situation (iii) in Thm. 4.1 where the refined retraining process is applied on the advantaged group $i$ while the original process is applied on the disadvantaged group $j$. From Fig. 19(a), 19(c) and 19(e), we can see unfairness reaches a minimum in the middle of the retraining process, suggesting the earlier stopping of retraining brings benefits. From Fig. 19(b), 19(d) and 19(f), we can see $\Delta_{t}$ for the disadvantaged group $j$ reaches a minimum in the middle of the retraining process, but generally not at the same time when unfairness reaches a minimum.
@@ F.6 Comparisons between Strategic and Non-strategic Situations
F.7 Additional Results on a Dataset with Non-Linearity.
ACSIncome-CA dataset [69] is a larger dataset consisting of over $150K$ records for agents and their annual income. The decision-maker wants to predict whether a person has an annual income $>50000$. We assume the decision-maker first trains a 2-d embedding using a neural network and $53$ original features then regards the embedding as the new feature. We divide the agents into 2 groups based on their ages. Similar to the credit approval dataset, we then fit Beta distributions on the 2 groups and then verify the monotonic likelihood assumption (Fig. 13). We then plot the dynamics of $a_{t},q_{t},\Delta_{t}$ for both groups when the systematic bias is either positive or negative. The results show that similar trends still hold for this large dataset (Fig.  14).
Figure 13: Feature distribution of Income Dataset.
Figure 14: Dynamics of $a_{t},q_{t},\Delta_{t}$ of the income dataset: the left plot is for group $i$ and the right plot is for group $j$. We run $10$ trials for each experiment and agent cost matrices are the same as the experiments in the main paper. The results are similar to the ones in the main paper.
@@ G.2 Derivation of factors influencing $a_{t},q_{t}$
According to the proof of Theorem 3.3, all agents with feature vector $x$ who are admitted by $f_{0}$ will be admitted by $f_{t}$ ($t>0$), making all agents who possibly improve must belong to $\mathcal{J}$. When $F_{X}$ is convex in $\mathcal{J}$ or it is convex in each dimension separately, we know $P_{X}$ as its derivative will be non-decreasing in each of the $d$ dimensions within $\mathcal{J}$. Similarly, when $P_{Y|X}(1|x)$ is convex in each of its dimensions or the whole $\mathcal{J}$, $U$ is also non-decreasing in each of the $d$ dimensions within $\mathcal{J}$. Then note that $f_{t}$ always lies below $f_{t-1}$, therefore, for each agent who improves at $t$ having feature vector $x_{i,t}$, we can find an agent at $t-1$ with corresponding $x_{i,t-1}$ such that both $P_{X}(x_{i,t-1})\geq P_{X}(x_{i,t})$ and $U(x_{i,t-1})\geq U(x_{i,t})$. This will ensure $I(t)\leq I(t-1)$. Thus, $q_{t}$ decreases starting from 1. ∎
@@
At round $t$, we know $A_{\theta_{t}^{i}}^{it}>A_{\theta_{t}^{j}}^{jt}$. To reach demographic parity, the acceptance rates need to be the same, so at least one of the following situations must happen: (i) $\widetilde{\theta}_{t}^{i}>\theta_{t}^{i}$ and $\widetilde{\theta}_{t}^{j}>\theta_{t}^{j}$; (ii) $\widetilde{\theta}_{t}^{i}<\theta_{t}^{i}$ and $\widetilde{\theta}_{t}^{j}<\theta_{t}^{j}$; (iii) $\widetilde{\theta}_{t}^{i}\geq\theta_{t}^{i}$ and $\widetilde{\theta}_{t}^{j}\leq\theta_{t}^{j}$. Next we prove that (i) and (ii) cannot be true by contradiction.
@@
Note that we want to sacrifice the least accuracy to reach demographic parity. Since the agent best response will never make them worse off, we will know $a_{t}^{i}$ corresponding to the acceptance rate of group $i$ under the optimal threshold $\theta_{t}^{i}$ is always larger than or equal to $a_{0}^{i}$. Thus, it suffices to show that under the condition of Thm. 4.2, $a_{t}^{j}<a_{0}^{i}$.
@@ Appendix H All plots with error bars
Although all our theoretical results are expressed in terms of expectation, we provide error bars for all plots in the main paper, App. E and F if randomness is applicable. All the above figures demonstrate expectations as well as error bars ($\pm$ 1 standard deviation). Overall, the experiments have reasonable standard errors. However, experiments in the Credit Approval dataset [20] (Fig. 36) incur larger standard errors, which is not surprising because the dataset violates several assumptions. Finally, note that we conduct 50-100 randomized trials for every experiment and we should expect much lower standard errors if the numbers of trials become large.